OpenStudy (anonymous):

WILL MEDAL AND FAN!!! Please Help!!!! Jenny is eight years older than twice her cousin Sue’s age. The sum of their ages is less than 32. What is the greatest age that Sue could be? The greatest age that Sue could be is ____ years old.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Thanks @MREAR53

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Jenny = x+8 Jenny + Sue <32 31-8 = 23 23/2=11.5 Sue is 11.5 or 11 Because the sum of their ages has to be less than 32